OpenStudy (henryblah):

Random thing I was thinking about. For example, when solving quadratics for example, I equal it to 0. x^2+2x-8=0 You can take out a factor of (x+4) to get (x+4)(x-2)=0. Then you can equate them both to 0 to get the answer. What I want to know is what is stopping me from taking out a random factor i.e. (x+7) (x+7)((x^2+2x-8)/(x+7))=0 And then getting x=-7. That's a simplified analogy and I know you wouldn't do it, but it gets more complicated when taking things out in trig.

OpenStudy (anonymous):

Naw, not really. I mean it's zero. Either way it doesn't really matter.

OpenStudy (henryblah):

? I'm not understanding what you mean.

OpenStudy (cwrw238):

x+ 7 / x + 7 = 0/0 if x = -7. 0/0 is meaningless

OpenStudy (kinggeorge):

In essence, if you do that, you'll be dividing by 0. And you can't do that.

OpenStudy (anonymous):

@cwrw238 said what I was trying to say. :)

OpenStudy (henryblah):

Ok. So if I had 4(cosx)^2-2-sinx-cosx=0 Could I divide by cosx and equate that to 0? i.e. cosx(4cosx)-2secx-tanx-1=0

OpenStudy (henryblah):

Whoops cosx(4cosx-2secx-tanx-1)=0 Brackets over entire thing

OpenStudy (kinggeorge):

I don't think that would work.

OpenStudy (henryblah):

Is there a reason why -.-?